# UGG-ReID: Uncertainty-Guided Graph Model for Multi-Modal Object Re-Identification

### v2

@@ 5 Conclusion
In this paper, we propose a novel Uncertainty-Guided Graph model for multi-modal object (UGG-ReID), effectively mining modal-specific information and boosting modal collaboration for multi-modal object ReID. This UGG-ReID consists of two main aspects. One design is the Gaussian Patch-Graph Representation (GPGR) model for quantifying the epistemic uncertainty of local and global features and modeling the relationship between them. This manner mitigates the noise interference and enhances the learning ability and robustness of modality-specific information. The other introduces the Uncertainty-Guided Mixture of Experts (UGMoE) strategy to select appropriate experts based on sample uncertainty and facilitate deep multi-modal interactions through a novel routing mechanism. These two core schemes synergistically optimize the overall network structure and achieve the current optimal performance on all five multi-modal object ReID datasets.
@@ A.1 Details of UGMoE
In the main text, we set the value of $k$ in Top $k$ to be 1. In this case, each modality will have $C-1$ unique experts and $M$ shared experts, so the total number of experts is $C+M-1$. If $k\neq 1$, we analyze it further to get $C+k(M-1)$ experts for each modality. Thus, Eq. 10 is defined as follows,
@@
The learned feature tends to be obtained by specific experts, which means that the existence of some experts can not be optimized. To solve this problem, we further add regular terms as,
@@
A.2 Details of Experiments
A.2.1 Experiments Setting
Datasets. To comprehensively evaluate the generalization ability of the proposed UGG-ReID framework, we conduct experiments on five public datasets. These include two person re-identification datasets, RGBNT201 [9] and Market1501-MM [36], as well as three challenging vehicle re-identification benchmarks: MSVR310 [10], RGBNT100 [8], and WMVEID863 [11]. These datasets collectively reflect a wide range of real-world scenarios and associated challenges. Table 4 summarizes the partition protocols and the specific challenges posed by each dataset.
@@ A.2.1 Experiments Setting
RGBNT201 Market1501-MM MSVR310 RGBNT100 WMVEID863
@@
A.2.2 Ablation Analysis
To verify the role of each loss in the model, we conduct systematic ablation experiments, as shown in Table 5. $\mathcal{L}_{c,s}$ represents the sum of $\mathcal{L}_{c}$ and $\mathcal{L}_{s}$, which is used to impose constraints on the global token. From the experimental results, one can observe that when the $\mathcal{L}_{c,s}$ constraint on the global token is removed, the performance of the model on multiple evaluation indicators decreases, indicating that the constraint has a positive effect on improving modeling ability. Then, $\mathcal{L}^{m}_{r}$ is removed to verify performance for adding the loss constraint on the expert. We can find that adding the loss, our mAP/R-1 increases by 2.2%/3.1% and 2.0%/1.6% in RGBNT201 [9] and WMVEID863 [11], respectively, which verifies its enhancement effect on the expert selection strategy Finally, we verify the effectiveness of $\mathcal{L}^{m}_{e}$, which aims to ensure that the number of similar samples assigned to each expert in the training process is balanced. Meanwhile, the expert weights are relatively evenly distributed among the experts, and the experimental results show that it can effectively prevent the imbalance of distribution among experts and improve the ability of the model.
A.2.3 Hyperparameter Analysis
We analyze the effects of the hyperparameters $C$ and $k$ on the model performance, where $C$ controls the number of experts, and $k$ denotes the number of share-experts selected for each modality. As shown in Table 6, a moderate increase in $C$ enhances the model’s expressive capacity, while an appropriate choice of $k$ strikes a balance between stability and flexibility. This facilitates dynamic collaboration and complementarity among experts, ultimately improving overall model performance.
@@ A.2.4 Visual Results
A.3 Discussion
Multi-modal object ReID exploits fine-grained local cues and the complementary information of modalities to effectively enhance the robustness and accuracy of recognition in complex scenarios [9, 20, 22, 27, 23]. As is well known, significant distributional differences exist among different modalities, and noise arising from sample quality and environmental factors further impacts the accuracy of feature representations. The proposed UGG-ReID effectively guides the feature fusion process by explicitly quantifying local and sample-level epistemic uncertainties and modeling the relationship between them, enhancing the model’s robustness and effectiveness. UGG-ReID is the first work that leverages uncertainty to quantify fine-grained-local details and explicitly model their dependencies in multi-modal data. In future work, we will focus on advancing uncertainty quantification and reasoning techniques, exploring the integration of Bayesian inference and evidence theory into multi-modal object ReID [41, 42, 43]. This aims to enhance the model’s robustness to modality and label noise, thereby improving its overall performance and reliability in complex environments.

### v3

@@ 5 Conclusion
In this paper, we propose a novel Uncertainty-Guided Graph model for multi-modal object (UGG-ReID), effectively mining modal-specific information and boosting modal collaboration for multi-modal object ReID. This UGG-ReID consists of two main aspects. One design is the Gaussian Patch-Graph Representation (GPGR) model for quantifying the aleatoric uncertainty of local and global features and modeling the relationship between them. This manner mitigates the noise interference and enhances the learning ability and robustness of modality-specific information. The other introduces the Uncertainty-Guided Mixture of Experts (UGMoE) strategy to select appropriate experts based on sample uncertainty and facilitate deep multi-modal interactions through a novel routing mechanism. These two core schemes synergistically optimize the overall network structure and achieve the current optimal performance on all five multi-modal object ReID datasets.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62372003, 62576004), the Natural Science Foundation of Anhui Province (2308085Y40, 2408085J037), the Key Technologies R&D Program of Anhui Province (202423k09020039), and the Open Research Project of the Anhui Provincial Key Laboratory of Security Artificial Intelligence (SAI202401).
@@ A.1 Details of UGMoE
In the main text, we set the value of $k$ in Top $k$ to be 1. In this case, each modality will have C1C-1 unique experts and $M$ shared experts, so the total number of experts is C+M1C+M-1. If $k\neq 1$, we analyze it further to get C+k(M1)C+k(M-1) experts for each modality. Thus, Eq. 10 is defined as follows,
@@
The learned feature tends to be obtained by specific experts, which means that the existence of some experts can not be optimized. To solve this problem, we further add regular terms [10.1145/3664647.3680949] as,
@@
A.2 Comparison with Prior Works
For conceptual comparison, we first utilize a Gaussian-based random graph for object representation, where nodes are described by Gaussian distributions to represent the uncertainty of image patches in the presence of noise, whereas previous works generally employ a deterministic graph model, represented by a feature vector. We design a Gaussian Patch-Graph Representation (GPGR) to quantify aleatoric uncertainties for global and local features while modeling their relationships. To our knowledge, this work is the first attempt to exploit a random patch-graph model for the object ReID problem. Second, for the Mixture-of-Experts approach, we design the Uncertainty-Guided Mixture of Experts (UGMoE) strategy, which enables different samples to select experts based on uncertainty and utilizes an uncertainty-guided routing mechanism to strengthen the interaction between multi-modal features, effectively promoting modal collaboration.
Table 6: Component-wise comparison of different methods on RGBNT201 (in %).
Method Uncer. MoE Local Gloabl Graph mAP R-1
EUAR [10.1145/3664647.3680949] $\times$ $\times$ 74.1 77.6
EAU [10657706] $\checkmark$ $\times$ $\times$ $\times$ 75.6 80.3
MAP [10205473] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ 76.8 78.2
DeMo [wang2024demo] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 79.7 81.8
UGG-ReID $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.2 86.8
For a technical comparison, we first further analyze the impact of different uncertainty modeling approaches on the performance of multi-modal object ReID [10657706, 10205473, 10.1145/3664647.3680949]. As shown in Table 6, works EUAR [10.1145/3664647.3680949] and EAU [10657706] are able to perceive inter-sample uncertainty. In contrast, MAP [10205473] introduces a more comprehensive uncertainty modeling mechanism to quantify the uncertainty of local cues. Although the above methods take uncertainty into account, they either neglect the modeling of uncertainty in local cues or the structural relationships between local regions. Second, we perform comparisons under different MoE strategies. As shown in Table 6, we substitute two existing MoE methods [10.1145/3664647.3680949, wang2024demo]. Compared with DeMo [wang2024demo], UGMoE better exploits the diversity of samples by introducing uncertainty modeling, and compared with EUAR [10.1145/3664647.3680949], UGMoE further strengthens the interaction between different modalities.
A.3 Details of Experiments
A.3.1 Experiments Setting
Datasets. To comprehensively evaluate the generalization ability of the proposed UGG-ReID framework, we conduct experiments on five public datasets. These include two person re-identification datasets, RGBNT201 [Zheng2021RobustMP] and Market1501-MM [Wang2022InteractEA], as well as three challenging vehicle re-identification benchmarks: MSVR310 [ZHENG2023101901], RGBNT100 [Li2020MultiSpectralVR], and WMVEID863 [ZHENG2025102800]. These datasets collectively reflect a wide range of real-world scenarios and associated challenges. Table 7 summarizes the partition protocols and the specific challenges posed by each dataset.
@@ A.3.1 Experiments Setting
RGBNT201 Market1501-$M$ MSVR310 RGBNT100 WMVEID863
@@
A.3.2 Ablation Analysis
To verify the role of each loss in the model, we conduct systematic ablation experiments, as shown in Table 8. $\mathcal{L}_{c,s}$ represents the sum of $\mathcal{L}_{c}$ and $\mathcal{L}_{s}$, which is used to impose constraints on the global token. From the experimental results, one can observe that when the $\mathcal{L}_{c,s}$ constraint on the global token is removed, the performance of the model on multiple evaluation indicators decreases, indicating that the constraint has a positive effect on improving modeling ability. Then, $\mathcal{L}^{m}_{r}$ is removed to verify performance for adding the loss constraint on the expert. We can find that adding the loss, our mAP/R-1 increases by 2.2%/3.1% and 2.0%/1.6% in RGBNT201 [Zheng2021RobustMP] and WMVEID863 [ZHENG2025102800], respectively, which verifies its enhancement effect on the expert selection strategy Finally, we verify the effectiveness of $\mathcal{L}^{m}_{e}$, which aims to ensure that the number of similar samples assigned to each expert in the training process is balanced. Meanwhile, the expert weights are relatively evenly distributed among the experts, and the experimental results show that it can effectively prevent the imbalance of distribution among experts and improve the ability of the model.
A.3.3 Hyperparameter Analysis
We analyze the effects of the hyperparameters $C$ and $k$ on model performance, where $C$ controls the number of experts and $k$ denotes the number of shared experts selected for each modality. As shown in Table 9, a moderate increase in $C$ enhances the model’s expressive capacity, while an appropriate choice of $k$ strikes a balance between stability and flexibility. This facilitates dynamic collaboration and complementarity among experts, ultimately improving overall model performance.
We further analyze the local nodes $n$ of GPGL and Layers $L$ of GPGCN of the GPGR for the effect of the model in the RGBNT201 dataset in Table 10. For nodes n, we observe that n=128 achieves excellent results. Too few nodes are not enough to cover rich local information, and too many introduce redundancy and noise, interfering with graph structure learning. For layers $L$, GPGCN works best when LL=2. The number of layers is too shallow and may lead to insufficient fusion of local structures, while too deep may cause over-smoothing, resulting in the loss of discriminative representation of nodes and weakening the expression ability of local discriminative features.
@@ A.3.3 Hyperparameter Analysis
Table 10: Results of the analysis on hyperparameters $L$ and $n$ on the RGBNT201 dataset (in %).
Layers $L$ RGBNT201
mAP R-1 mAP R-1
1 80.1 85.3 91.3 93.9
2 81.2 86.8 92.0 94.7
3 79.1 84.6 90.8 93.3
4 78.4 82.5 90.3 92.8
5 76.3 80.7 89.1 91.7
Nodes $n$ RGBNT201
mAP R-1 mAP R-1
32 79.3 83.7 91.3 93.9
64 80.1 82.9 90.0 94.0
96 81.4 84.6 90.4 92.6
128 81.2 86.8 92.0 94.7
160 79.8 83.6 90.2 91.9
@@ A.3.4 Visual Results
A.4 Discussion
Multi-modal object ReID exploits fine-grained local cues and the complementary information of modalities to effectively enhance the robustness and accuracy of recognition in complex scenarios [Zheng2021RobustMP, Wang2023TOPReIDMO, wang2025IDEA, Zhang25PromptMA, wang2024demo]. As is well known, significant distributional differences exist among different modalities, and noise arising from sample quality and environmental factors further impacts the accuracy of feature representations. The proposed UGG-ReID effectively guides the feature fusion process by explicitly quantifying local and sample-level epistemic uncertainties and modeling the relationship between them, enhancing the model’s robustness and effectiveness. UGG-ReID is the first work that leverages uncertainty to quantify fine-grained-local details and explicitly model their dependencies in multi-modal data.
Limitations and In the Future. Our framework employs uncertainty-guided learning to enhance robustness against local noise; it may still struggle under extreme conditions where local cues are heavily corrupted or missing. In future work, we will focus on advancing uncertainty quantification and reasoning techniques, exploring the integration of Bayesian inference and evidence theory into multi-modal object ReID [Kaushik_2024_CVPR, DCEL, UGNCL]. This aims to enhance the model’s robustness to modality and label noise, thereby improving its overall performance and reliability in complex environments.